OpenStudy (anonymous):

The area of the largest rectangle that can be drawn with one side along the x-axis and two vertices on the curve of y=e^-x^2 is...

OpenStudy (dumbcow):

so length of base of rectangle is 2x height of rectangle is y or e^(-x^2) this gives an area function:

OpenStudy (dumbcow):

\[A = 2x e^{-x^2}\] differentiate and set equal to zero to maximize area

OpenStudy (anonymous):

ok so derivative is \[A'= 2e ^{-x ^{2}}\]

OpenStudy (anonymous):

or do you have to take derivative of -x^2 also?

OpenStudy (dumbcow):

using product rule \[\frac{dA}{dx} = 2 e^{-x^2} -4x^2 e^{-x^2} = 0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what so we do next?

OpenStudy (dumbcow):

wait do you understand product rule and how i got derivative?

OpenStudy (anonymous):

yea, i just always forget to use it :/

OpenStudy (anonymous):

f'(x)= a'b +ab'

OpenStudy (dumbcow):

yeah ok, next just solve for x

OpenStudy (dumbcow):

factor out the e^-x^2 part

OpenStudy (anonymous):

so e^-x^2(2-4x^2)

OpenStudy (anonymous):

=0

OpenStudy (dumbcow):

yes e^-x^2 = 0 has no solution solve 2-4x^2 = 0

OpenStudy (anonymous):

x=sqrt 1/2

OpenStudy (anonymous):

so base of rectangle is 1

OpenStudy (dumbcow):

hmm no , base = 2x ----> 2*sqrt(1/2)

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

:/ so base is 2rt1/2 and then height is e^-1/2?

OpenStudy (dumbcow):

you can rationalize sqrt(1/2) to sqrt2/2

OpenStudy (dumbcow):

yep so max Area = 2sqrt(1/2)*e^-1/2 or sqrt(2)*e^-1/2

OpenStudy (anonymous):

ok, thank you a lot! I understand the concept, but i was making some silly math errors.

OpenStudy (dumbcow):

yw :)